OpenStudy (anonymous):

Help with algebra 2?

OpenStudy (anonymous):

Sure, what is the question?

OpenStudy (michele_laino):

I think that your answer is correct!

OpenStudy (michele_laino):

@crustymcbeth

OpenStudy (anonymous):

assume :- u= x b, then x=u1b  and dx=1b du x b−1 ∫ ∞ 0 x a 1+ x b = 1 b ∫ ∞ 0 x a (1+u) x b−1 du = 1 b ∫ ∞ 0 x a−b+1 (1+u) du = 1 b ∫ ∞ 0 u a+1 b −1 (1+u) du  let m= a+1 b  and n+m=1 then n=1− a+1 b 1 b ∫ ∞ 0 u a+1 b −1 (1+u) du = 1 b ∫ ∞ 0 u m−1 (1+u ) n+m du = 1 b β(m,n) = 1 b β( a+1 b ,1− a+1 b ) = 1 b Γ( a+1 b )Γ(1− a+1 b ) Γ( a+1 b +1− a+1 b ) = 1 b Γ( a+1 b )Γ(1− a+1 b ) Γ(1) =1b π sin(π(a+1b)) we have to go in the complex plane,namely I set: x (b/2)=z so your integral becomes: 2b ∫ 0 ∞ z 2a/b+2/b−1 1+ z2dz now that integral is equal to: 2π/2 (Resf(i)+Resf(−i)) where Res is the residual value of f(z) and f(z) is our integrand above. after a standard calculation, we get: πb4 (e) iπ(a+1)/b−e −iπ(a+1)/b) =πb4 ∗2isin(πa+1b) and taking the real part we can write: πb2 ∗sin(πa+1b)

OpenStudy (michele_laino):

please you have to use this properties of logarithms: \[n*\log _{5}x=\log _{5}(x)^{n}\] where n can be 0, 1, 2, 3,... or 0, -1, -2, -3,...as well \[\log _{5}x+\log _{5}y=\log _{5}(x+y)\] and this: \[\frac{ 1 }{ n }\log _{5}x=\log _{5}\sqrt[n]{x}\] where n, can be equal to 1,2,3,4,...

OpenStudy (michele_laino):

oops... these properties...

OpenStudy (michele_laino):

furthermore, this property can be useful: \[\log _{5}x-\log _{5}y=\log _{5}\left( \frac{ x }{ y } \right)\]

OpenStudy (michele_laino):

please, try to apply those properties above, in order to simplify your expression

OpenStudy (anonymous):

i still dont know what to do

OpenStudy (solomonzelman):

let me briefly post the rules: \(\Large\color{royalblue}{ c\cdot\log_a(b)= \log_a (b^c) }\) (exponent) \(\Large\color{blueviolet}{\log_a(d)+\log_a(f)= \log_a (d \cdot f) }\) (addition) \(\Large\color{royalblue}{\log_a(d)-\log_a(f)= \log_a (d \div f) }\) (subtraction)

OpenStudy (anonymous):

u have to know the formulas

OpenStudy (anonymous):

@SolomonZelman has mentioned them

OpenStudy (solomonzelman):

apply the rules respectively.... and you should be alright